# Design and synthesis of scalable quantum programs

### v1

@@ 1 Introduction
Since 2017, we have seen massive and continuous advances in quantum hardware in many modalities, including superconducting devices [7, 8, 9, 10], ion traps [11, 12, 13], neutral atoms [14, 15], silicon devices [16, 17], photonic devices [18, 19], and more [20]. The consensus is that quantum computing has shifted from a scientific/academic phase into an engineering phase during this time frame. The different hardware roadmaps vary, but all show major applicative breakthroughs that are expected in the next few years [21, 22, 23, 24, 25, 26, 27, 28]. Software concepts and platforms have not evolved at the same pace. The first software platforms [4, 29, 30, 31] appeared hand-in-hand with the public exposition of quantum hardware accessible on the cloud in 2016 [32]. These platforms typically consisted of a means to specify the operations of quantum gates (and later quantum pulses) on a set of qubits, transpilation of the gate-level programs into equivalent gate-level programs suitable to a given hardware, and software for simulating the quantum hardware on classical CPUs and GPUs. Advancements in those software offerings included high-level constructs such as quantum registers and algorithmic building blocks. However, those were not true abstractions of the problem because the constructs only encapsulated hard-coded, low-level implementations of qubits and quantum gates; these were packaged as named-functions performing operations on named registers. Further developments in quantum software included application libraries built on top of the basic software, including applications in quantum machine learning, finance, chemistry, and many more [4, 5, 33, 34]. These libraries provided further domain-specific algorithms and functions, based on building blocks tuned and optimized for a specific application. Fundamentally, these were not different from hard-coded quantum circuits encapsulated in several levels of building block functions. In addition, several quantum programming languages that describe functionality at a higher level of abstraction in different respects have been proposed in recent years [3, 35, 36]. The tool chains for these languages utilize compilation frameworks and techniques adapted from classical programming language processing, notably MLIR [37] and LLVM [38]. The actual compilation approach in all these cases still rigidly maps the abstractions onto a low-level implementation, taking only local resource considerations into account. We present a radically different approach to the compilation of quantum programs. Our strategy follows and extends the general ideas presented earlier [39, 40, 2]. However, for the first time the ideas are fully implemented and do not remain at the conceptual levels. The level of understanding and detail gained through implementation clearly changes and cements the conceptual level as well; hence the work presented here creates a major step forward in advancing the path to true, scalable quantum software. At the base of our work is abstraction. By abstraction we mean allowing the user to express any information of interest to them regarding the quantum program, and omit any other information, or detail, that is not of interest. It is up to the technology to fill in all details that were not expressly specified, and come up with an executable, optimal, quantum program that conforms to the specification of the user and fits a given hardware. The abstract description from the user is a model of the program, while the outcome executable code is a concrete instance of the model. In this paper, we will focus on functional models for which the abstract description describes required functionality of the program, constraints on the hardware, and possible optimization criteria. At this level of description, there are many details that need to be filled. For example, each function may have many types of implementation, and each type can be implemented in itself in many different ways. An adder for example may be implemented as a simple ripple-carry adder, or as a Fourier transformed adder in which computation is done in quantum Fourier space. The result of the adder can be put out of place on a newly allocated register of qubits, leaving the two input registers intact, or can replace one or both of the input variables. Each operation within the adder may use any number of auxiliary qubits (’auxiliaries’) in order to perform the operation faster or with fewer two-qubit gates. Finally, the adder may be implemented to any level of accuracy, again affecting the number of qubits and two-qubit gates that the operation will consume. In all those cases, the functionality prescribed by the user remains intact [^1] Playing with the accuracy degree of freedom results in different functionalities of the adder, however we consider a likely model in which only the accuracy of the entire program is specified, and not the accuracy of each component.. Beyond the single function, unspecified details may also relate to the interaction between functions. For example, different functions may or may not share auxiliary or functional qubits, depending on the scope of the function and whether its results are required downstream, and also on the cost of disentangling and resetting used qubits. More generally, the decision on which qubits, if at all, to disentangle and reuse, as well as the ordering of functions using the same auxiliary qubits, is a very important ’detail’ from the implementation point of view, but is of no concern to the user from a functional model point of view. Such considerations of ’filling in the details’ between wide levels of abstraction are the basis of many electronic design automation (EDA) technologies, specifically synthesis from high-level functional and nonfunctional models to register-transfer-level (RTL) models and then to logic, netlist, and physical-level implementations. Given the circuit model of quantum computation, it should not be surprising that this general synthesis approach may also be applicable in the quantum domain. In this paper, we present the application of the EDA synthesis approach to quantum computing. Throughout the paper, we use the word synthesis as its meaning in the EDA industry, which is much more general than the term sometimes used in quantum computing [41] - meaning the decomposition of a specific unitary into a given gate set. Our results below demonstrate the viability of this approach for creating robust quantum software at scale. When developing the technology, we tried to preserve as much as possible all the guidelines and best practices of the conventional principles of EDA synthesis. The challenges we encountered in applying those principles to the nature of quantum programs were already significant enough so that we did not attempt to invent any new wheel when it was not necessary. Still, the peculiarities of quantum programs, including no fan-in or fan-out, no copy, disentanglement as means for freeing qubit resources, long-range correlations, inability to locally simulate sub-circuits, and many more – make the incorporation of these concepts into EDA synthesis principles non-trivial and challenging. In this paper, we mainly focus on the synthesis technology aspect per se. While we present modeling examples when needed, we do not delve into the research behind the definition of the quantum modeling constructs and the common quantum idioms that we found leading to those constructs. In addition, we do not describe here the challenges behind analysis and debugging of the automatically synthesized quantum program, and feedback into improving the model from this analysis. The modeling constructs and automatic program analysis are important parts of the EDA flow that we have again adapted into the quantum software cycle but are beyond the scope of this paper.
@@ 2.1 Quantum walk on a circle
Fig. 2 also shows a comparison to the best-of-the-art model implementations using several other quantum software tools. In all those cases, implementations have a fixed CX-count for any number of control qubits in the function: exponential number in Qiskit [4] and Pennylane [5], and a cubic number in TKet [29]). We note that the libraries of these tools do have ad hoc, more efficient implementations for the MCX operation in terms of the CX gate count. However, when using these more efficient implementations, auxiliary qubits must be declared and treated as part of the function itself. Implementing the program with the compilers of these tools, which lack auxiliary awareness and management, results in a large circuit width because the auxiliaries of each function add up instead of being reused. In general terms, the underlying concept of all existing standard tools (including those that do include memory management [54, 35, 55]) is that they do not distinguish between the notion of a function and that of an implementation.
@@
Figure 2: Large scale quantum walk programs on a circle of $2^{N}$ nodes. The classiq curves are obtained by synthesizing with optimization over the CX counts and constraining the total width to 40 (blue dots) or 100 (brown stars). (a,b) The total CX counts, and the total generation time respectively (we stop at compilation times of $O(10^{3})$sec). The inset in panel (b) indicates the auxiliary usage. For comparison, the results for a similar implementation using various quantum platforms are plotted as well (see Methods for technical detail, and code for generating the results in Ref. [56]).
@@ 2.2 Quantum Singular Value Transform
$A=\begin{pmatrix}0&0&0&\dots&0&0\\ 0.5&0&0.5&\dots&0&0\\ 0&0.5&0&0.5&\dots&0\\ \vdots&0&\ddots&0&\dots&0\\ \vdots&0&\ddots&0.5&0&0.5\\ \end{pmatrix}.$ (1)
@@
Figure 3: Block encoding model for the matrix $A$ in Eq. (1). (a): A schematic description of the model, the $I-|0\rangle\langle 0|$ function is defined with single-qubit gates and an MCX operation, see Methods. (b,c): Two different implementations for $N=20$ obtained by the synthesis engine when optimizing over CX-counts and constraining the width to 45 (b) or 90 (c) qubits. Different adders are chosen (QFT or ripple carry, respectively), given the different constraints, resulting in CX-counts of 1480 and 842 for a 44 and 47 qubits circuit in panels (b) and (c), respectively.
Given a block encoding model, one can plug it into a QSVT routine to block encode a polynomial of degree $l$, $P_{l}(A)$ [57]. The model corresponds to applications of $U_{A}$, its inverse, and rotations about the block states; see Methods. We generate a random odd polynomial of degree 3 and apply it to different values of the matrix size, see Fig. 4. Our optimized results for CX counts, compared to state-of-the-art results from other software tools, originates from the flexibility of choosing efficient implementations of controlled operations and quantum adders.
Figure 4: QSVT for a polynomial of matrix $A$ in Eq. (1) for varying matrix sizes $2^{N}$. (a,b) Total CX counts and total generation time, where the inset of panel (b) presents the auxiliary counts. The Classiq curves are obtained by optimizing over the CX counts and constraining the total width to 50 (brown stars) or 100 (blue dots).
@@ 3 Conclusions
We have discussed the application of EDA synthesis methods to the field of quantum software and have shown clear results that support this approach as the robust, scalable and long-lasting way to approach quantum software development. One thing we have not covered in detail are the affect of different optimization functions and in particular fitness to a given hardware. For example, the graphs of Figs. 2,4 show results for which the synthesis engine optimizes for a minimal number of CX gates. They would look very different if the user decided to optimize for minimal width, providing much narrower circuits to the user. In addition, taking target hardware into account may have a crucial effect on the synthesis. For example, in fully connected hardware the synthesis engine may be more likely to implement an MCX gate with more CX gates than when the target is sparsely connected hardware, because in the latter case the cost of each CX gate is very large because of the need for SWAPS. We stress that such decisions on how to implement an MCX or other high-level functions cannot be taken by a transpiler which accepts a low-level implementation as input. Only true abstraction and true dynamic synthesis, as in this paper, can make such crucial decisions. Lastly, we stress that the decision on how to implement a function may be different for each instance of the function in the program, depending on the resources available locally for the function at this part of the program.
@@ 4 Methods
A detailed description of the synthesis engine is provided in the Supplementary Materials. Here we provide some technical details needed to perform the experiments of Sec. 2, while the corresponding code is available in Ref. [56].
@@ 4.1 Numerical study details
In all platforms, we use the tool’s compilation and transpilation abilities to get a final quantum program with only one-qubit gate operations and CX operations. This allows a fair comparison for the CX counts per example. All runtime experiments were performed on a single Apple M1 pro, all on the same computer with the same software installed and the same environmental conditions. During run-time, the software in question was the only significant application running on the computer.
@@ 4.2 Quantum walk model
The models in all other platforms are designed in the same fashion, using the platforms’ building blocks such as quantum control operation, inversion, and MCX gates [56].
@@ 4.3 QSVT on the block-encoded $A$
We block encode the matrix $A$ in Eq. (1) as a product of $A_{0}=0.5\sum^{2^{N}-2}_{i=0}(|i\rangle\langle i+1|+0.5|2^{N}-1\rangle\langle 0|+c.c$ and $A_{1}=I-|0\rangle\langle 0|$, , and apply block-encoding multiplication. The block encoding of $A_{0}$ is done with 2 block qubits, according to the procedure described in [63] for Toeplitz matrices [see Eq. (55) therein]. The matrix $A_{1}$ can be block encoded as a Linear Combination of Unitaries (LCU) for the unitaries $I$ and $I-2|0\rangle\langle 0|$. The latter unitary is the reflection around zero operation appearing in Grover’s Search algorithm. The Qmod code for block encoding $A$ is given in Listing 3.

### v2

@@ 1 Introduction
Since 2017, we have seen massive and continuous advances in quantum hardware in many modalities, including superconducting devices [7, 8, 9, 10], ion traps [11, 12, 13], neutral atoms [14, 15], silicon devices [16, 17], photonic devices [18, 19], and more [20]. The consensus is that quantum computing has shifted from a scientific/academic phase into an engineering phase during this time frame. The different hardware roadmaps vary, but all show major applicative breakthroughs that are expected in the next few years [21, 22, 23, 24, 25, 26, 27, 28].
Software concepts and platforms have not evolved at the same pace. The first software platforms [4, 29, 30, 31] appeared hand-in-hand with the public exposition of quantum hardware accessible on the cloud in 2016 [32]. These platforms typically consisted of a means to specify the operations of quantum gates (and later quantum pulses) on a set of qubits, transpilation of the gate-level programs into equivalent gate-level programs suitable to a given hardware, and software for simulating the quantum hardware on classical CPUs and GPUs. Advancements in those software offerings included high-level constructs such as quantum registers and algorithmic building blocks. However, those were not true abstractions of the problem because the constructs only encapsulated hard-coded, low-level implementations of qubits and quantum gates; these were packaged as named-functions performing operations on named registers.
Further developments in quantum software included application libraries built on top of the basic software, including applications in quantum machine learning, finance, chemistry, and many more [4, 5, 33, 34]. These libraries provided further domain-specific algorithms and functions, based on building blocks tuned and optimized for a specific application. Fundamentally, these were not different from hard-coded quantum circuits encapsulated in several levels of building block functions. In addition, several quantum programming languages that describe functionality at a higher level of abstraction in different respects have been proposed in recent years [3, 35, 36]. The tool chains for these languages utilize compilation frameworks and techniques adapted from classical programming language processing, notably MLIR [37] and LLVM [38]. The actual compilation approach in all these cases still rigidly maps the abstractions onto a low-level implementation, taking only local resource considerations into account.
We present a radically different approach to the compilation of quantum programs. Our strategy follows and extends the general ideas presented earlier [39, 40, 2]. However, for the first time the ideas are fully implemented and do not remain at the conceptual levels. The level of understanding and detail gained through implementation clearly changes and cements the conceptual level as well; hence the work presented here creates a major step forward in advancing the path to true, scalable quantum software.
At the base of our work is abstraction. By abstraction we mean allowing the user to express any information of interest to them regarding the quantum program, and omit any other information, or detail, that is not of interest. It is up to the technology to fill in all details that were not expressly specified, and come up with an executable, optimal, quantum program that conforms to the specification of the user and fits a given hardware. The abstract description from the user is a model of the program, while the outcome executable code is a concrete instance of the model. In this paper, we will focus on functional models for which the abstract description describes required functionality of the program, constraints on the hardware, and possible optimization criteria.
At this level of description, there are many details that need to be filled. For example, each function may have many types of implementation, and each type can be implemented in itself in many different ways. An adder for example may be implemented as a simple ripple-carry adder, or as a Fourier transformed adder in which computation is done in quantum Fourier space. The result of the adder can be put out of place on a newly allocated register of qubits, leaving the two input registers intact, or can replace one or both of the input variables. Each operation within the adder may use any number of auxiliary qubits (’auxiliaries’) in order to perform the operation faster or with fewer two-qubit gates. Finally, the adder may be implemented to any level of accuracy, again affecting the number of qubits and two-qubit gates that the operation will consume. In all those cases, the functionality prescribed by the user remains intact [^1] Playing with the accuracy degree of freedom results in different functionalities of the adder, however we consider a likely model in which only the accuracy of the entire program is specified, and not the accuracy of each component..
Beyond the single function, unspecified details may also relate to the interaction between functions. For example, different functions may or may not share auxiliary or functional qubits, depending on the scope of the function and whether its results are required downstream, and also on the cost of disentangling and resetting used qubits. More generally, the decision on which qubits, if at all, to disentangle and reuse, as well as the ordering of functions using the same auxiliary qubits, is a very important ’detail’ from the implementation point of view, but is of no concern to the user from a functional model point of view.
Such considerations of ’filling in the details’ between wide levels of abstraction are the basis of many electronic design automation (EDA) technologies, specifically synthesis from high-level functional and nonfunctional models to register-transfer-level (RTL) models and then to logic, netlist, and physical-level implementations. Given the circuit model of quantum computation, it should not be surprising that this general synthesis approach may also be applicable in the quantum domain. In this paper, we present the application of the EDA synthesis approach to quantum computing. Throughout the paper, we use the word synthesis as its meaning in the EDA industry, which is much more general than the term sometimes used in quantum computing [41] - meaning the decomposition of a specific unitary into a given gate set. Our results below demonstrate the viability of this approach for creating robust quantum software at scale.
When developing the technology, we tried to preserve as much as possible all the guidelines and best practices of the conventional principles of EDA synthesis. The challenges we encountered in applying those principles to the nature of quantum programs were already significant enough so that we did not attempt to invent any new wheel when it was not necessary. Still, the peculiarities of quantum programs, including no fan-in or fan-out, no copy, disentanglement as means for freeing qubit resources, long-range correlations, inability to locally simulate sub-circuits, and many more – make the incorporation of these concepts into EDA synthesis principles non-trivial and challenging.
In this paper, we mainly focus on the synthesis technology aspect per se. While we present modeling examples when needed, we do not delve into the research behind the definition of the quantum modeling constructs and the common quantum idioms that we found leading to those constructs. In addition, we do not describe here the challenges behind analysis and debugging of the automatically synthesized quantum program, and feedback into improving the model from this analysis. The modeling constructs and automatic program analysis are important parts of the EDA flow that we have again adapted into the quantum software cycle but are beyond the scope of this paper.
@@ 2.1 Quantum walk on a circle
Fig. 2 also shows a comparison to the best-of-the-art model implementations using several other quantum software tools. In all those cases, implementations have a fixed CX-count for any number of control qubits in the function, which scales quadratically with the number of control variable [^3] In TKet [29] this results in a cubic trend in the main plot of Fig. 2. The steeper curves of Qiskit [4] and Pennylane [5] for the CX-counts probably originates in inefficient compilation of the controlled-MCX function, not interpreting it as a larger MCX gate.. We note that the libraries of these tools do have ad hoc, more efficient implementations for the MCX operation in terms of the CX gate count. However, when using these more efficient implementations, auxiliary qubits must be declared and treated as part of the function itself. Implementing the program with the compilers of these tools, which lack auxiliary awareness and management, results in a large circuit width because the auxiliaries of each function add up instead of being reused. In general terms, the underlying concept of all existing standard tools (including those that do include memory management [54, 35, 55]) is that they do not distinguish between the notion of a function and that of an implementation.
@@
Figure 2: Large scale quantum walk programs on a circle of $2^{N}$ nodes. The Classiq curves are obtained by synthesizing with optimization over the CX counts and constraining the total width to 40 (blue dots) or 100 (brown stars). The main plot shows the total CX counts, and the two insets (a,b) shows the total generation time and programs’ width, respectively (we stop at compilation times of $\sim 10^{3}$ sec).
@@ 2.2 Quantum Singular Value Transform
$A=\begin{pmatrix}0&0&0&\dots&0&0\\ 0.5&0&0.5&\dots&0&0\\ 0&0.5&0&0.5&\dots&0\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&0&0.5&0&0.5\\ 0&0&0&0&0.5&0\end{pmatrix}.$ (1)
@@
Figure 3: Block encoding model for the matrix $A$ in Eq. (1). (a): A schematic description of the model, the $I-2|0\rangle\langle 0|$ function is defined with single-qubit gates and an MCX operation, see Methods. (b,c): Two different implementations for $N=20$ obtained by the synthesis engine when optimizing over CX-counts and constraining the width to 45 (b) or 90 (c) qubits. Different adders are chosen, QFT or ripple carry (rca), given the different constraints, resulting in CX-counts of 1480 and 842 for a 44 and 47 qubits circuit in panels (b) and (c), respectively.
Given a block encoding model, one can plug it into a QSVT routine to block encode a polynomial of degree $l$, $P_{l}(A)$ [56]. The model corresponds to applications of $U_{A}$, its inverse, and rotations about the block states; see Methods. We generate a random odd polynomial of degree 3 and apply it to different values of the matrix size. Fig. 4. shows our optimized results for CX counts, compared to state-of-the-art results from other software tools, originates from the flexibility of choosing efficient implementations of controlled operations and quantum adders.
Figure 4: QSVT for a polynomial of matrix $A$ in Eq. (1) for varying matrix sizes $2^{N}$. The main plot shows the total CX counts, and the two insets (a,b) shows and total generation time and the quantum program width respectively (we stop at compilation times of $\sim 10^{3}$ sec). The Classiq curves are obtained by optimizing over the CX counts and constraining the total width to 50 (blue dots) or 100 (brown stars).
@@ 3 Conclusions
We have discussed the application of EDA synthesis methods to the field of quantum software and have shown clear results that support this approach as the robust, scalable and long-lasting way to approach quantum software development. One thing we have not covered in detail are the affect of different optimization functions and in particular fitness to a given hardware. For example, the graphs of Figs. 2 and 4 show results for which the synthesis engine optimizes for a minimal number of CX gates. They would look very different if the user decided to optimize for minimal width, providing much narrower circuits to the user. In addition, taking target hardware into account may have a crucial effect on the synthesis. For example, in fully connected hardware the synthesis engine may be more likely to implement an MCX gate with more CX gates than when the target is sparsely connected hardware, because in the latter case the cost of each CX gate is very large due to the need of swap gates for routing. We stress that such decisions on how to implement an MCX or other high-level functions cannot be taken by a transpiler which accepts a low-level implementation as input. Only true abstraction and true dynamic synthesis, as we employ in this paper, can make such crucial decisions. Lastly, we stress that the decision on how to implement a function may be different for each instance of the function in the program, depending on the resources available locally for the function at this part of the program.
@@ 4 Methods
A detailed description of the synthesis engine is provided in the Supplementary Materials. Here we provide some technical details needed to perform the experiments of Sec. 2, while the corresponding code is available in a public repository [62].
@@ 4.1 Numerical study details
In all platforms, we use the tool’s compilation and transpilation abilities to get a final quantum program with only one-qubit gate and CX gate operations. This allows a fair comparison for the CX counts per example. All runtime experiments were performed on a single Apple M1 pro, all on the same computer with the same software installed and the same environmental conditions. During run-time, the software in question was the only significant application running on the computer.
@@ 4.2 Quantum walk model
The models in all other platforms are designed in the same fashion, using the platforms’ building blocks such as quantum control and inversion operations [62].
@@ 4.3 QSVT on the block-encoded $A$
We block encode the matrix $A$ in Eq. (1) as a product of $A_{0}=0.5\sum^{2^{N}-2}_{i=0}(|i\rangle\langle i+1|+0.5|2^{N}-1\rangle\langle 0|+c.c$ and $A_{1}=I-|0\rangle\langle 0|$, and apply block-encoding multiplication. The block encoding of $A_{0}$ is done with 2 block qubits, according to the procedure described in [63] for Toeplitz matrices [see Eq. (55) therein]. The matrix $A_{1}$ can be block encoded as a Linear Combination of Unitaries (LCU) for the unitaries $I$ and $I-2|0\rangle\langle 0|$. The latter unitary is the reflection around zero operation appearing in Grover’s Search algorithm. The Qmod code for block encoding $A$ is given in Listing 3.
@@
Conflict of interest/Competing interests
The authors declare no competing interests.
Data and Code availability
The datasets generated during and/or analyzed during the current study, and the code that generates them, are partially available in Ref. [62]. The rest of the code is available from the corresponding author on reasonable request.